Nanticoke Industries had the following operating results for 2018: sales $29,500; cost of goods sold = $19,760; depreciation expense = $5,260; interest expense = $2,640; dividends paid = $1,450. At the beginning of the year, net fixed assets were $17,080, current assets were $5,860, and current liabilities were $3,325. At the end of the year, net fixed assets were $20,660, current assets were $7,300, and current liabilities were $3,980. The tax rate for 2018 was 35 %. a. What is net income for 2018? (Negative answers should be indicated by a minus sign. Omit $ sign in your response.) Net income $ b. What is the operating cash flow for 2018? (Negative answers should be indicated by a minus sign. Omit $ sign in your response.) Operating cash flow $ c. What is the cash flow from assets for 2018? (Negative answer should be indicated by a minus sign. Omit $ sign in your response.) Cash flow from assets $ d. If no new debt was issued during the year, what is the cash flow to creditors? What is the cash flow to shareholders? (Negative answers should be indicated by a minus sign. Omit $ sign in your response.) Cash flow to creditors $ Cash flow to shareholders $
Answer #1

SOLUTION

a.

Sales 29,500
Cost of goods sold 19,760
Gorss profit 9,740
Depreciation Expenses 5,260
Interest 2,640
Operating Income 1,840
Income Tax @35% 644
Net income 1,196

b. Operating Cash flows for 2018 = Gross Profit - Income Tax

= 9,740 - 644 = 9,096

c. Beginning net working capital = Beginning Current Assets - Beginning Current Liabilities

= 5,860 - 3,325 = 2,535

Ending net working capital = Ending Current Assets - Ending Current Liabilities

= 7,300 - 3,980 = 3,320

Change in net working capital = Ending net working capital - Beginning net working capital

= 3,320 - 2,535 = 785

Net Capital Spending = Ending Net Fixed Assets - Beginning Net Fixed Assets + Depreciation

= 20,660 - 17,080 + 5,260 = 8,840

Cash flow from Assets = Operating Cash flow - Change in net working capital - Net capital spending

= 9,096 - 785 - 8,840 = (529)

d1. Cash flow to Creditor = Interest Expense - Net new Long-term Debt

= $2,640 - 0 = $2,640

d2. Cash flow from Assets = Cash flow to Creditor + Cash flow to Stockholders

-529 = 2,640 + Cash flow to Stockholders

Cash flow to Stockholders = 2,640 + 529 = 3,169